\medskip

\begin{thebibliography}{9}

\bibitem{Bra85} L. De Branges, \emph{A proof of the Bieberbach conjecture}, Acta Math. {\bf 154} (1985), no. 1--2, 137--152.


\bibitem{duren}
P.L. Duren, Univalent function, Springer-Verlag, New York, 1983.

\bibitem{FekSze33} M. Fekete and G. Szeg\"o, \emph{Eine Bemerkung \"uber ungerade schlichte Funktionen}, J. London Math. Soc. {\bf 8} (1933), 85--89.

\bibitem{Gri76} A. Z. Grinspan, \emph{The sharpening of the difference of the moduli of adjacent coefficients of schlicht functions}, In Some problems in modern function theory Proc. Conf. Modern Problems of Geometric Theory of Functions, Inst. Math., Acad. Sci. USSR, Novosibirsk, (Russian), Akad. Nauk SSSR Sibirsk. Otdel. Inst. Mat., Novosibirsk, (1976), 41--45.



\bibitem{Hay63} W. K. Hayman, \emph{On successive coefficients of univalent functions}, J. London Math. Soc., {\bf 38} (1963), 228--243.



\bibitem{Lebedev}
N.A. Lebedev, Area principle in the theory of univalent functions, Published by"Nauka", Moscow, 1975 (in Russian).

\bibitem{OT} Milutin Obradovi\'{c} and Nikola Tuneski, \emph{Hankel determinants of second and third  order for the class of univalent functions}, (submitted).


\bibitem{DTV-book}
D.K. Thomas, N. Tuneski, A. Vasudevarao, Univalent Functions: A Primer, \emph{De Gruyter Studies in Mathematics} {\bf 69}, De Gruyter, Berlin, Boston, 2018.



\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb}
\usepackage{graphicx}

\newtheorem{thm}{Theorem}
\newtheorem{cor}{Corollary}
\newtheorem{lem}{Lemma}

\newtheorem{rem}{Remark}

\newcommand{\es}{{\mathcal S}}

\newcommand{\D}{{\mathbb D}}

\def\be{\begin{equation}}
\def\ee{\end{equation}}

\newcommand{\bee}{\begin{enumerate}}
\newcommand{\eee}{\end{enumerate}}

\newcommand{\blem}{\begin{lem}}
\newcommand{\elem}{\end{lem}}
\newcommand{\bthm}{\begin{thm}}
\newcommand{\ethm}{\end{thm}}
\newcommand{\bcor}{\begin{cor}}
\newcommand{\ecor}{\end{cor}}
\newcommand{\beg}{\begin{example}}
\newcommand{\eeg}{\end{example}}
\newcommand{\begs}{\begin{examples}}
\newcommand{\eegs}{\end{examples}}
\newcommand{\bdefe}{\begin{defin}}
\newcommand{\edefe}{\end{defin}}
\newcommand{\bprob}{\begin{prob}}
\newcommand{\eprob}{\end{prob}}
\newcommand{\bei}{\begin{itemize}}
\newcommand{\eei}{\end{itemize}}

\newcommand{\bcon}{\begin{conj}}
\newcommand{\econ}{\end{conj}}
\newcommand{\bcons}{\begin{conjs}}
\newcommand{\econs}{\end{conjs}}
\newcommand{\bprop}{\begin{propo}}
\newcommand{\eprop}{\end{propo}}
\newcommand{\br}{\begin{rem}}
\newcommand{\er}{\end{rem}}
\newcommand{\brs}{\begin{rems}}
\newcommand{\ers}{\end{rems}}
\newcommand{\bo}{\begin{obser}}
\newcommand{\eo}{\end{obser}}
\newcommand{\bos}{\begin{obsers}}
\newcommand{\eos}{\end{obsers}}
\newcommand{\bpf}{\begin{pf}}
\newcommand{\epf}{\end{pf}}
\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\beq}{\begin{eqnarray}}
\newcommand{\beqq}{\begin{eqnarray*}}
\newcommand{\eeq}{\end{eqnarray}}
\newcommand{\eeqq}{\end{eqnarray*}}

\begin{document}
\bibliographystyle{amsplain}

\title[On the difference of coefficients of univalent functions]{On the difference of coefficients of univalent functions}

\author[M. Obradovi\'{c}]{Milutin Obradovi\'{c}}
\address{Department of Mathematics,
Faculty of Civil Engineering, University of Belgrade,
Bulevar Kralja Aleksandra 73, 11000, Belgrade, Serbia}
\email{obrad@grf.bg.ac.rs}

\author[D. K. Thomas]{Derek K. Thomas}
\address{Department of Mathematics, Swansea University, Bay Campus, Swansea, SA1 8EN, United Kingdom}
\email{d.k.thomas@swansea.ac.uk}

\author[N. Tuneski]{Nikola Tuneski}
\address{Department of Mathematics and Informatics, Faculty of Mechanical Engineering, Ss. Cyril and Methodius
University in Skopje, Karpo\v{s} II b.b., 1000 Skopje, Republic of North Macedonia.}
\email{nikola.tuneski@mf.edu.mk}

\subjclass[2000]{30C45, 30C50, 30C55}
\keywords{Univalent Functions, Grunsky Coefficients, Coefficient Differences }




\begin{abstract}
For $f\in \mathcal{S}$, the class of normalized functions, analytic and univalent in the unit disk $\mathbb{D}$ and given by $f(z)=z+\sum_{n=2}^{\infty} a_n z^n$ for $z\in \mathbb{D}$, we give an upper bound for the coefficient difference $|a_4|-|a_3|$  when $f\in \mathcal{S}$. This  provides an improved bound in the case $n=3$  of Grispan's 1976 general bound $||a_{n+1}|-|a_n||\le 3.61\dots .$ Other coefficients bounds, and bounds for the second and third Hankel determinants when $f\in \mathcal{S}$ are found when either $a_2=0,$ or $a_3=0$.
\end{abstract}

\maketitle

\medskip

\section{Introduction. preliminaries and definitions}

\medskip

Let $\mathcal{A}$ be the class of functions $f$ which are analytic  in the open unit disc $\D=\{z:|z|<1\}$ of the form
\be\label{e1}
f(z)=z+a_2z^2+a_3z^3+\cdots,
\ee
and let $\mathcal{S}$ be the subclass of $\mathcal{A}$ consisting of functions that are univalent in $\D$.

\medskip

Although the famous Bieberbach conjecture $|a_n|\le n$ for $n\ge2,$ was proved by de Branges  in 1985 \cite{Bra85}, a great many other problems concerning the coefficients $a_n$ remain open. The main aim of this paper (Section 3), is by use of  the Grunsky inequalities, to find an upper for the difference of coefficients $|a_4|-|a_3|$ for $f\in \mathcal{
S}$, which improves the well-known general bound of Grispan $||a_{n+1}|-|a_n||\le 3.61\dots$  \cite{Gri76}, when $n=3$.
We also obtain information concerning the initial coefficients of $f(z)$, and of the second and third Hankel determinants  when either $a_2=0$, or $a_3=0$.
\bigskip

For $f\in \mathcal{S}$, the Grunsky coefficients $\omega_{p,q}$ as defined in N. A. Lebedev \cite{Lebedev} are given by
\[  \log\frac{f(t)-f(z)}{t-z}=\sum_{p,q=0}^{\infty}\omega_{p,q}t^{p}z^{q},  \]

\medskip

\noindent where  $\omega_{p,q}=\omega_{q,p}$, and satisfy the so-called Grunsky inequalities \cite{duren,Lebedev}

\be\label{eq 4}
\sum_{q=1}^{\infty}q \left|\sum_{p=1}^{\infty}\omega_{p,q}x_{p}\right|^{2}\leq \sum_{p=1}^{\infty}\frac{|x_{p}|^{2}}{p},
\ee
where $x_{p}$ are arbitrary complex numbers such that last series converges.

\medskip

Further, it is well-known that if $f$ given by \eqref{e1}
belongs to $\mathcal{S}$, then also
\be\label{eq 5-mo-4}
f_{2}(z)=\sqrt{f(z^{2})}=z +c_{3}z^3+c_{5}z^{5}+\cdots
\ee
belongs to $\mathcal{S}$. Thus for the function $f_{2}$ we have the appropriate Grunsky
coefficients of the form $\omega_{2p-1,2q-1}$, and  inequalities \eqref{eq 4} take the form
\be\label{eq 6-mo-5}
\sum_{q=1}^{\infty}(2q-1) \left|\sum_{p=1}^{\infty}\omega_{2p-1,2q-1}x_{2p-1}\right|^{2}\leq \sum_{p=1}^{\infty}\frac{|x_{2p-1}|^{2}}{2p-1}.
\ee

\medskip

(Note that in this paper, we omit the upper index (2) in  $\omega_{2p-1,2q-1}^{(2)}$ in Lebedev's notation).

\medskip

The following similar inequality follows from the relation (15) on page 57 in \cite{Lebedev}.
\be\label{eq 6-mo-6}
 \left|\sum_{p=1}^{\infty}\sum_{q=1}^{\infty}\omega_{2p-1,2q-1}x_{2p-1}x_{2q-1}\right|\leq \sum_{p=1}^{\infty}\frac{|x_{2p-1}|^{2}}{2p-1}.
\ee

\medskip

Thus for example, from  \eqref{eq 6-mo-5} and \eqref{eq 6-mo-6} when $x_{2p-1}=0$ and $p=3,4,\ldots$, we obtain
\begin{equation}\label{e7}
|\omega_{11} x_1 +\omega_{31} x_3 |^2 +3|\omega_{13} x_1 +\omega_{33} x_3 |^2 + 5|\omega_{15} x_1 +\omega_{35} x_3 |^2 \le |x_1|^2+\frac{|x_3|^2}{3}
\end{equation}
and
\begin{equation}\label{e8}
|\omega_{11} x_1^2 +2\omega_{13}x_1 x_3 +\omega_{33}x_3^2 |\le |x_1|^2+\frac{|x_3|^2}{3},
\end{equation}
respectively.

\medskip

It was also shown in \cite[p.57]{Lebedev}, that if $f\in\mathcal{S}$ is given by \eqref{e1}, then the coefficients $a_{2}$, $ a_{3}$, $ a_{4}$ and $a_5$ can be expressed in terms of the Grunsky coefficients  $\omega_{2p-1,2q-1}$ of the function $f_{2}$ given by
\eqref{eq 5-mo-4} as follows.
\be\label{e9}
\begin{split}
a_{2}&=2\omega _{11},\\
a_{3}&=2\omega_{13}+3\omega_{11}^{2}, \\
a_{4}&=2\omega_{33}+8\omega_{11}\omega_{13}+\frac{10}{3}\omega_{11}^{3},\\
a_{5}&=2\omega_{35}+8\omega_{11}\omega_{33}+5\omega_{13}^{2}+18\omega_{11}^2\omega_{13}+\frac73\omega_{11}^4,\\
0&= 3\omega_{15}-3\omega_{11}\omega_{13}+\omega_{11}^3-3\omega_{33}.
\end{split}
\ee
In this paper we will use these expressions to obtain information concerning the coefficients $a_2$, $a_3$, $a_4$, and $a_5$ when $f\in\mathcal{S}$.
\medskip

In recent years a great deal of attention has been given to finding upper bounds for the modulus of the second and third Hankel determinants $H_2(2)$ and $H_3(1)$, defined as follows who's elements are the coefficients of $f\in\mathcal{S}$ (see e.g. \cite{DTV-book}).
 \medskip

For $f\in\mathcal{S}$
\[  H_2(2) = a_2a_4-a_3^2  \]
and
\begin{equation}\label{e11}
H_3(1) = a_3(a_2a_4-a_3^2) - a_4(a_4-a_2a_3)+a_5(a_3-a_2^2).
\end{equation}
\medskip

Almost all results have concentrated on finding bounds for $|H_2(2)|$ and $|H_3(1)|$ for subclasses of $\mathcal{S}$, and only recently has a significant bound been found for the whole class $\mathcal{S}$ \cite{OT} for $|H_2(2)|$ and $|H_3(1)|$. However finding exact sharp bounds remains an open problem.
\medskip

We begin by using the Grunsky inequalities in (\ref{eq 6-mo-6}) to obtain bounds for the modulus of some initial coefficients and $|H_2(2)|$ and $|H_3(1)|$ when  $f\in \mathcal{S}$ provided either $a_2,$ or $a_3=0.$

\medskip
\section{Coefficient bounds and Hankel determinants}

\medskip

Obtaining sharp bounds for the modulus of the coefficients for odd functions in $\mathcal{S}$ has long been been an open problem. If $f_2$, given by (\ref{eq 5-mo-4}) is an odd function in $\mathcal{S}$, then  the only known sharp bounds for $|c_{2n-1}|$ for $n\ge 2$ are $|c_3|\le 1$, and $|c_5|\le 1/2+e^{-2/3}=1.013\dots$. In general the best bound to date is $|c_{2n-1}|\le1.14 $ for $n\ge2$, (see e.g.\cite{duren}).
\medskip

In our first theorem, we give  bounds for $|a_3|$, $|a_4|$ and $|a_5|$ when $f\in \mathcal{S}$   assuming only that only $a_2=0,$ thus providing bounds for a wider class of functions than the odd functions in $\mathcal{S}$. We also give bounds for $|H_2(2)|$ and $|H_3(1)|$  in this case.
\begin{thm}\label{th-1}
Let $f\in\es$ and be given by \eqref{e1} with $a_2=0$. Then
\begin{itemize}
\item[($i$)] $|a_3|\le1$,
\medskip
\item[($ii$)] $|a_4|\le\frac23=0.666\ldots$,
\medskip
\item[($iii$)]  $|a_5|\le \sqrt{\frac{19}{15}} = 1.67666\ldots$,
\medskip
\item[($iv$)] $|H_2(2)|\le1$,
\medskip
\item[($v$)] $|H_3(1)|\le\frac{21}{20} = 1.05$.
\end{itemize}
\end{thm}
\medskip
\begin{proof}$\mbox{}$
\begin{itemize}
\item[($i$)] The classical inequality $|a_3-a_2^2|\le1$ for $f$ in $\es$ when $a_2=0$, gives  $|a_3|\le1$, which from (\ref{e9}) gives
\be\label{e12}
|\omega_{13}|\le\frac12.
\ee
\medskip
\item[($ii$)] Next choose  $x_1=0$ and $x_3=1$ in \eqref{e8}, which gives
\be\label{e13}
|\omega_{33}| \le\frac13.
\ee
Also, since $\omega_{11}=0$ ($\Leftrightarrow a_2=0$), then from \eqref{e9} and \eqref{e13} we obtain
\[
|a_4|=2|\omega_{33}|\le\frac23=0.666\ldots.
\]
\medskip
\item[($iii$)]  Again since  $\omega_{11}=0$,  from \eqref{e9} we obtain
\be\label{e14}
|a_5|=|2\omega_{35}+5\omega_{13}^2|.
\ee

From \eqref{e7} with $x_1=0$ and $x_3=1$ we have ($\omega_{11}=0$)
\[  |\omega_{13}|^2 + 3|\omega_{33}|^2 + 5|\omega_{35}|^2 \le \frac13 \]
and from here
\begin{equation}\label{n1}
  |\omega_{35}| \le \frac{1}{\sqrt{15}} \sqrt{1-3|\omega_{13}|^2}.
\end{equation}
From \eqref{e14} and \eqref{n1} we have
\[ |a_{5}| \le 2|\omega_{35}|+5|\omega_{13}|^2 \le \frac{1}{\sqrt{15}} \sqrt{1-3|\omega_{13}|^2}+5|\omega_{13}|^2 \le \frac{503}{300}=1.67666\ldots.\]

\medskip

\item[($iv$)] Since we are assuming  $a_2=0$, $(i)$ shows that   $|H_2(2)| \le1$ is trivial.
\medskip

\item[($v$)] When $\omega_{11}=0$, from the last relation in \eqref{e9} we have $\omega_{33}=\omega_{15},$ and from \eqref{e11},
\be\label{e16}
\begin{split}
|H_3(1)
&= |2\omega_{13}^3+4\omega_{13}\omega_{35}-4\omega_{33}^2|\le 2|\omega_{13}|^3 + 4 + \underbrace{|\omega_{13}\omega_{35} -\omega_{15}^2|}_{E_1}.
\end{split}
\ee

Now choose $x_1=-\omega_{15}$, and $x_3=\omega_{13}$, and since $\omega_{33}=\omega_{15}$, from \eqref{e7} we obtain
\[ |\omega_{13}|^4+5E_1^2\le |\omega_{15}|^2+\frac{|\omega_{13}|^2}{3} \le \frac15-\frac35|\omega_{13}|^2  +\frac13|\omega_{13}|^2,\]
(since by \eqref{e7} $3|\omega_{13}|^2+5|\omega_{15}|^2\le1$ for $x_1=1$, $x_3=0$ and $\omega_{11}=0$),
which implies $5E_1^2\le \frac15-\frac{4}{15}|\omega_{13}|^2-|\omega_{13}|^4$, i.e., $E_1\le \frac15$.

\medskip

Finally from \eqref{e12} and \eqref{e16}, it follows that
\[ |H_3(1)| \le 2\cdot \frac18+4\cdot \frac15 = \frac{21}{20}=1.05. \]
\end{itemize}
This completes the proof of Theorem \ref{th-1}.
\end{proof}

\medskip

We next prove a similar result, this time assuming that $a_3=0$.
\begin{thm}\label{th-2}
Let $f\in\es$ and be  given by \eqref{e1}, with $a_3=0$. Then
\medskip
\begin{itemize}
\item[($i$)] $|a_2|\le1$,
\medskip
\item[($ii$)] $|a_4|\le \frac{\sqrt{37}+13}{12}=1.59023\ldots$,
\medskip
\item[($iii$)]  $|a_5|\le \frac14 \sqrt{\frac{757}{15}}+ \frac{85}{64} = 3.10412\ldots$,
\medskip
\item[($iv$)] $|H_2(2)|\le\frac{13+\sqrt{37}}{12}=1.59023\ldots$,
\medskip
\item[($v$)] $|H_3(1)|\le \frac{24+\sqrt{645}}{30}  = 1.64656\ldots$.
\end{itemize}
\end{thm}
\medskip

\begin{proof}$\mbox{}$
\begin{itemize}
\item[($i$)] Since $|a_3-a_2^2|\le1$ and $a_3=0$, then $|a_2^2|\le1$, i.e., $|a_2|\le1$. Also, since by \eqref{e9}, $a_3=2\omega_{13}+3\omega_{11}^2=0$, it follows that
\be\label{e17}
\omega_{13}=-\frac32 \omega_{11}^2 \quad \left(\Leftrightarrow\,\, \omega_{11}^2=-\frac23\omega_{13}\right).
\ee
Because $|a_2|=|2\omega_{11}|\le1$, we have
\be\label{e18}
|\omega_{11}|\le \frac12 \quad\mbox{and}\quad |\omega_{13}|\le \frac38 \,\,(\mbox{by } \eqref{e17}.
\ee
\medskip
\item[($ii$)] By using \eqref{e9} and \eqref{e17}, we obtain
\be\label{e19}
\begin{split}
|a_4|
&= \left|2\omega_{33}+8\omega_{11} \left(-\frac32\omega_{11}^2\right)+\frac{10}{3}\omega_{11}^3\right|\\
&= \left|2\omega_{33}-\frac{26}{3}\omega_{11}^3\right| \\
&\le 2|\omega_{33}|+\frac{26}{3}|\omega_{11}| ^3.
\end{split}
\ee
From \eqref{e7}, using $x_1=0$ and $x_3=1$, we have
\[  |\omega_{13}|^2 + 3|\omega_{33}|^2 \le \frac13, \]
which implies (with $\omega_{13}=-\frac32 \omega_{11}^2$, see \eqref{e17})
\be\label{e20}
|\omega_{33}|\le\sqrt{\frac19-\frac34 |\omega_{11}|^4}.
\ee
Combining \eqref{e19} and \eqref{e20} we obtain
\be\label{e21}
|a_4|\le 2 \sqrt{\frac19 -\frac34|\omega_{11}|^4} + \frac{26}{3}|\omega_{11}|^3 =: \varphi(|\omega_{11}|),
\ee
where $\varphi(t) =  2 \sqrt{\frac19 -\frac34 t^4} + \frac{26}{3}t^3 $, $0\le t =|\omega_{11}|\le\frac12$ (by \eqref{e18}). Since $\varphi$ is increasing function on  $[0,1/2]$,
\[ \varphi(t)\le \varphi(1/2)=\frac{\sqrt{37}+13}{12}, \]
which, together   with \eqref{e21}, gives the desired result.
\medskip
\item[($iii$)] From the last relation in \eqref{e9}, using \eqref{e17}  we have $\omega_{33}=\omega_{15}+\frac{11}{6}\omega_{11}^3$, which with the expression for $a_5$ in \eqref{e9}, gives
\be\label{e22}
\begin{split}
|a_5|
&= |2\omega_{35} + 8\omega_{11}\omega_{15} +5\omega_{13}^2 -10\omega_{11}^4|\\
&\le 2\underbrace{| \omega_{35} + 4\omega_{11}\omega_{15}|}_{C_1^\ast} + \underbrace{5|\omega_{13}|^2 +10|\omega_{11}|^4}_{C_2^\ast}.
\end{split}
\ee
Once again, using \eqref{e7} choosing $x_1=4\omega_{11}$, $x_3=1$ and $\omega_{13}=-\frac32 \omega_{11}^2$, we have
\[(C_1^\ast)^2 = |4\omega_{11}\omega_{15} +\omega_{35} |^2 \le -\frac54 |\omega_{11}|^4 +\frac{16}{5}|\omega_{11}|^2+\frac{1}{15} \le \frac{757}{64\cdot 15}, \]
since $|\omega_{11}|\le\frac12$. Thus
\[ C_1^\ast \le \frac18\sqrt{\frac{757}{15}}.\]

\medskip

Next, since $\omega_{13}=-\frac32 \omega_{11}^2$ and $|\omega_{11}|\le\frac12$, we have
\[
\begin{split}
C_2^\ast
&= 5\cdot\frac94\cdot |\omega_{11}|^4+10|\omega_{11}|^4 = \frac{85}{4}|\omega_{11}|^4 \le \frac{85}{4}\cdot \frac{1}{16} = \frac{85}{64},
\end{split}
\]
since $|\omega_{11}|\le\frac12$.

\medskip

Finally from  \eqref{e22} we have
\[  |a_5| \le \frac14 \sqrt{\frac{757}{15}}+ \frac{85}{64} = 3.10412\ldots.\]
\medskip
\item[($iv$)] By using \eqref{e11}, \eqref{e9} and \eqref{e17}, we have
\be\label{e24}
\begin{split}
H_2(2)
&= 4\omega_{11}\omega_{33} + 4\omega_{11}^2\omega_{13} - 4\omega_{13}^2 - \frac73\omega_{11}^4\\
&= 4\omega_{11}\omega_{33}-\frac{52}{3}\omega_{11}^4
\end{split}
\ee
and from here
\be\label{n2}
\begin{split}
|H_2(2)| \le 4|\omega_{11}||\omega_{33}|+\frac{52}{3}|\omega_{11}|^4.
\end{split}
\ee

From \eqref{e20} and \eqref{n2} we have
\[ |H_2(2)| \le 4|\omega_{11}|\sqrt{\frac19-\frac34|\omega_{11}|^4} +\frac{52}{3}|\omega_{11}|^4 =: \varphi_1(|\omega_{11}|,\]
where
\[ \varphi_1(t) = 4t\sqrt{\frac19-\frac34t^4}+\frac{52}{3}t^4, \]
with $0\le t=|\omega_{11}|\le\frac12$. Finally, it can be checked that  $\varphi_1$ is an increasing function on the interval $(0,1/2)$, and so
\[ |H_2(2)| \le \varphi_1(1/2) = \frac{13+\sqrt{37}}{12}=1.59023\ldots. \]

\medskip
\item[($v$)]
By using the last relation from \eqref{e9} with $\omega_{13}=-\frac32\omega_{11}^2$, it follows that $\omega_{33}=\omega_{15}+\frac{11}{6}\omega_{11}^3$, and so using  \eqref{e11}, after some calculations we obtain
\[ H_3(1) = -12\omega_{11}^2 \left(\omega_{11}\omega_{15} +\frac23\omega_{35}\right) - 4\omega_{15}^2 -30\omega_{11}^6,\]
which gives
\be\label{e25}
|H_3(1)| \le \underbrace{12 |\omega_{11}|^2 \left|\omega_{11}\omega_{15} +\frac23 \omega_{35}\right|}_{D_1} + \underbrace{4|\omega_{15}|^2  + 30|\omega_{11}|^6}_{D_2}.
\ee
Now choose $x_1=\omega_{11}$ and $x_3=\frac23$ in \eqref{e7}, then (since $\omega_{13}=-\frac32\omega_{11}^2$),
\[  \left|\omega_{11}\omega_{15} +\frac23\omega_{35}\right| \le \sqrt{\frac15\left( |\omega_{11}|^2+\frac{4}{27} \right)},\]
and so
\be\label{e26}
\begin{split}
D_1 &\le 12 |\omega_{11}|^2 \sqrt{\frac15\left(|\omega_{11}|^2+\frac{4}{27}\right)} \\
&\le 12 \cdot\frac14 \sqrt{\frac15\left(\frac14+\frac{4}{27}\right)} = \sqrt{\frac{43}{60}}=\frac{\sqrt{645}}{30}=0.84656\dots,
\end{split}
\ee
since $|\omega_{11}|\le\frac12$.

\medskip

Also, as in the proof of ($iii$), we have
\[ 5|\omega_{15}|^2\le 1-|\omega_{11}|^2 -3|\omega_{13}|^2 = 1-|\omega_{11}|^2-\frac{27}{4}|\omega_{11}|^4,  \]
where we have once again used $\omega_{13}=-\frac32\omega_{11}^2$.
Now
\[ D_2 \le \frac45 - \frac{4}{5}|\omega_{11}|^2 - \frac{27}{5}|\omega_{11}|^4 + 30|\omega_{11}|^6  =: \varphi_2(|\omega_{11}|^2), \]
where
\[  \varphi_2(t)=\frac15\left(4 -4t - 27t^2 +150t^3\right),  \]
and $0\le t = |\omega_{11}|^2 \le \frac14$. Since $\varphi_2$ attains its maximum at $t_0 = 0$,
\be\label{e29}
D_2 \le \varphi_2(0) = \frac45.
\ee
Finally, by using \eqref{e25}, \eqref{e26} and \eqref{e29} we obtain
\[ |H_3(1)| \le D_1+D_2 \le  \frac{24+\sqrt{645}}{30}  = 1.64656\ldots. \]
\end{itemize}
\end{proof}

\medskip

\section{Coefficient  differences for $f\in \mathcal{S}$}
\medskip

A long standing problem in the theory of univalent functions is to find  sharp upper and lower bounds for $|a_{n+1}|-|a_n|$, when $f\in\mathcal{S}$. Since  the Keobe function has coefficients $a_n=n$, it is natural to conjecture  that $||a_{n+1}|-|a_n||\le1$. As early as 1933, this was shown to be false even when $n=2,$ when Fekete and Szeg\"o \cite{FekSze33} obtained the sharp bounds
 $$ -1 \leq |a_3| - |a_2| \leq \frac{3}{4} + e^{-\lambda_0}(2e^{-\lambda_0}-1) = 1.029\ldots, $$
where $\lambda_0$ is the unique value of $\lambda$ in $0 < \lambda <1$, satisfying the equation $4\lambda = e^{\lambda}$.

\medskip

Hayman \cite{Hay63} showed that if $f \in {\mathcal S}$, then $| |a_{n+1}| - |a_n| | \leq C$, where $C$ is an absolute constant. The exact value of $C$ is unknown,  the best estimate to date
being $C=3.61\ldots$ \cite{Gri76}, which because of the sharp estimate above when $n=2$, cannot be reduced to $1$.

\medskip

We now use the methods of this paper to obtain a better upper bound in the case $n=3$.

\medskip

\begin{thm}
Let $f\in\es$ and be given by \eqref{e1}. Then
\[|a_4|-|a_3| \le 2.1033299\ldots.\]
\end{thm}

\medskip

\begin{proof}
 By using \eqref{e9} we have
\[|a_4|-|a_3| \le |a_4|-|\omega_{11}||a_3| \le  |a_4-\omega_{11} a_3| = 2\Big|\underbrace{\omega_{33} +3\omega_{11}\omega_{33} + \frac16 \omega_{11}^3}_{B}\Big|. \]
From \eqref{e8} with $x_1=\frac{1}{\sqrt6}\omega_{11}$ and $x_3=1$, we obtain
\[
\begin{split}
 & \left| \omega_{33} +\frac{2}{\sqrt6}\omega_{11}\omega_{13} +\frac16\omega_{11}^3 \right| \le \frac16 |\omega_{11}|^2 +\frac13  \\
 \Rightarrow \quad & \left|B+\left( \frac{2}{\sqrt6}-3 \right) \omega_{11}\omega_{13} \right| \le \frac16 |\omega_{11}|^2 +\frac13  \\
 \Rightarrow \quad & |B| \le \left(3- \frac{\sqrt6}{3} \right) |\omega_{11}||\omega_{13}| + \frac16 |\omega_{11}|^2 +\frac13  \\
 \Rightarrow \quad & |B| \le \left(3- \frac{\sqrt6}{3} \right) |\omega_{11}| \cdot \frac{1}{\sqrt3}\sqrt{1-|\omega_{11}|^2} + \frac16 |\omega_{11}|^2 +\frac13  \\
 \Rightarrow \quad & |B| \le \frac13\left[ (3\sqrt3-\sqrt2) |\omega_{11}| \sqrt{1-|\omega_{11}|^2} + \frac12 |\omega_{11}|^2 + 1 \right] =: \varphi(|\omega_{11}|),
\end{split}
\]
where $\varphi(t)= \frac13\left[ (3\sqrt3-\sqrt2) t \sqrt{1-t^2} + \frac12 t^2 + 1 \right]$ for $0\le t\le1$, and where we have used that $|\omega_{13}| \le \frac{1}{\sqrt3}\sqrt{1-|\omega_{11}|^2}$. Since the function $\varphi$ attains its maximum at
\[t_0 = \sqrt{\frac12+\frac16\sqrt{\frac{1}{379}(39+8\sqrt6)}} = 0.75202\ldots,\]
and since $\varphi(t_0) = \frac{1}{12}\left(5+\sqrt{117-24\sqrt6}\right)$, it follows that
\[ |a_4|-|a_3| \le 2\varphi(t_0) = 2.10495\ldots. \]
\end{proof}

\end{document}